A body of mass 2kg travels according to the law x(t)=pt+qt^2+rt^3 where p=3m/s q=4m/s^2 and r=5m/s^3.The force acting on the body at t=2s is

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
Force acting on a body is given by F=ma
orF = m\frac{\partial^2 x}{\partial t^2}
Now x=3t +4t2+5t
v = \frac{\mathrm{dx} }{\mathrm{d} t}
v= 3+8t+15t2
a = \frac{\mathrm{dv} }{\mathrm{d} t}
a=8+30t
Now a at 2 sec will be 68 
so force acting on the object will be ma =136N